OpenStudy (anonymous):

2ln 3x +1 = 5 Could you find the solution to this please?

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Subtract 1 both the sides..

OpenStudy (anonymous):

ok.. then what

OpenStudy (anonymous):

then divide by 2 both the sides..

OpenStudy (anonymous):

alrighty then?

OpenStudy (anonymous):

show me your work...

OpenStudy (anonymous):

ln3x=2?

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

good, now what?

OpenStudy (anonymous):

Take antilog both the sides: \[3x = Antilog(2)\]

OpenStudy (anonymous):

is the answer 7?

OpenStudy (anonymous):

sorry i meant 4

OpenStudy (anonymous):

What is antilog (2) i don't know..

OpenStudy (anonymous):

I dont know myself

OpenStudy (anonymous):

\[Antilog(2) = 7.39\] Divide it by 3 and you wll get x..

OpenStudy (anonymous):

What are the options ??

OpenStudy (anonymous):

2.46

OpenStudy (anonymous):

Yes I got the same ..

OpenStudy (anonymous):

Thank You so much, I'm almost done... log 3 5 + log 3 x = log 3 10 can you help me with this one?

OpenStudy (anonymous):

Use: \[\log(a) + \log(b) = \log(ab)\]

OpenStudy (anonymous):

\[\large \log_3(5x) = \log_3(10)\] Now logs are same so the terms inside will be equal: \[5x = 10\] find x from here..

OpenStudy (anonymous):

Nice now can you just check this one for me? 22n = 16-, i said the solution is -2

OpenStudy (anonymous):

22n = 16 ??

OpenStudy (anonymous):

22n = 16^-1 ....sorry

OpenStudy (anonymous):

no sorry again its 2^2n=16^-1

OpenStudy (anonymous):

\[\large 2^{2n} = 16^{-1} \implies 2^{2n} = 2^{-4}\] \[2n = -4\] \(n = -2\)

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

Thank you sooo much :)

OpenStudy (anonymous):

Welcome dear..